# The Probabilistic Foundations of Surveillance Failure: From False Alerts to Structural Bias

### v1

@@ 4.4 Numerical Example: Geographic Disparities
5 Bayesian Posterior Reliability and the Base-Rate Trap
5.1 The Bayesian Framework and Positive Predictive Value
We introduce standard notation from diagnostic testing and forensic statistics (Balding, 2002):
$\pi$ $\equiv\Pr(\text{target})\quad\text{(base rate)},$
By Bayes’ rule, the positive predictive value (PPV) is
$\Pr(\text{target}\mid\text{flag})=\frac{s\,\pi}{s\,\pi+q\,(1-\pi)}.$ (12)
$\mathbb{E}[\#\text{ flags}]\approx r\,s+n\,q.$
Thus, the expected false discovery rate (FDR) is
matching (12) exactly.
5.2 Posterior Reliability and Bayesian Critical Scales
Proposition 5.1 (Bayesian Critical Population for Actionable PPV).
$\Pr(\text{target}\mid\text{flag})\geq\alpha$
is satisfied whenever
$n\leq\frac{(1-\alpha)\,r\,s}{\alpha\,q}.$ (14)
Remark 5.2 (Regime of Validity).
Proposition 5.1 applies to the “needle in a haystack” setting where $r$ is fixed or grows slowly while $n$ increases. If instead $r\propto n$ (constant prevalence), then the ratio $r/n$ is fixed, and the condition for actionable PPV reduces to a bound on $q/s$ relative to $\alpha/(1-\alpha)$, independent of $n$. The sparse-target regime is the most challenging case for screening systems and is therefore the primary focus of this analysis.
Proof.
Remark 5.3 (The Bayesian Trap).
5.3 Likelihood Ratios and Classical Fallacies
$L=\frac{s}{q},$
and mistakenly infer that $\Pr(\text{target}\mid\text{flag})$ is therefore large. This is the classical prosecutor’s fallacy (Balding, 2002). Bayes’ rule shows:
5.4 Resolving the DNA Database Controversy
Forensic statisticians have long debated the evidential value of DNA “cold hits.” Stockmarr (1999) argued that searching databases weakens evidence by inflating coincidental match probabilities; Balding and Donnelly (1995); Balding (2002) countered that likelihood ratios preserve evidential weight. Our framework resolves the debate.
Using (16),
Standard STR genotype profiling yields match probabilities on the order of $q\sim 10^{-12}$ or smaller. Even with databases of size $n\sim 10^{6}$, the product $nq\sim 10^{-6}\ll 1$ remains far below the critical scale where coincidental matches become probable. In this extreme regime, likelihood ratios dominate the posterior odds, validating Balding’s argument that database size does not meaningfully dilute evidential weight.
By contrast, multi-attribute surveillance systems often operate far above the critical threshold, where prior odds collapse faster than likelihood ratios can compensate. In this regime, Stockmarr’s caution applies: match evidence loses evidential weight as search populations grow. The same mathematics governs both contexts; the numerical parameters place them on opposite sides of the critical scale $n\sim\exp(\lambda D)$.
5.5 Key Takeaways
  1. 1.
  2. 2.
  3. 3.
    Exponential data growth overwhelms adaptation: Reducing $q$ by a factor of $\beta$ requires increasing $k$ by only $\log(1/\beta)/(pD)$, while real-world data growth is exponential: $k(t)=k_{0}\gamma^{t}$ (Section 3). Thus posterior collapse is temporally inevitable.
  4. 4.
  5. 5.
Remark 5.4 (Connections to Classical Statistical Fallacies).
This section unifies the base-rate fallacy, the prosecutor’s fallacy, false discovery rate control (Benjamini and Hochberg, 1995), and the PPV problem in medical screening (Welch and Black, 2011). Conceptually, these phenomena are identical: all reflect Bayes’ rule under low prevalence and imperfect specificity. The widely cited argument that “most published research findings are false” (Ioannidis, 2005) is the same FDR/PPV problem in another domain.
6 Effective Dimensionality Under Correlation
Real surveillance data exhibit spatial and temporal dependencies that fundamentally alter tail probabilities. While correlation does not change the expected number of matches (by linearity of expectation, $\mathbb{E}[\sum_{i=1}^{k}X_{i}]=kp$ regardless of dependence structure), it causes overdispersion that inflates tail probabilities beyond the independent Poisson approximation.
6.1 Variance Inflation and Rate Function Reduction
For binary indicators $X_{1},\ldots,X_{k}$ with $\mathbb{E}[X_{i}]=p$ and pairwise correlation $\text{Corr}(X_{i},X_{j})=\rho_{ij}$, the sum $Y=\sum_{i=1}^{k}X_{i}$ has:
$\mathbb{E}[Y]$ $=kp\quad\text{(unchanged by correlation)},$
$\text{Var}(Y)$ $=kp(1-p)+\sum_{i\neq j}p(1-p)\rho_{ij}=kp(1-p)\left(1+\frac{\sum_{i\neq j}\rho_{ij}}{k}\right),$
where the design effect $\text{DEFF}=\frac{\sum_{i\neq j}\rho_{ij}}{k}$ quantifies variance inflation; for nonnegative average correlation this satisfies $\text{DEFF}\geq 0$ (Kish, 1965).
For the Poisson approximation, independence gives $\text{Var}(Y)=\lambda=kp$. Under positive correlation, $\text{Var}(Y)>\lambda$, creating heavier tails. Positive correlation generally reduces the large-deviation rate compared to the independent case, producing heavier tails; see Dembo and Zeitouni (2010) for conditions. For intuition, one can treat the reduction in rate as shrinking the exponent by the factor $k_{\text{eff}}/k$.
6.2 Effective Degrees of Freedom
Rather than reducing $\lambda$ itself, correlation reduces the effective degrees of freedom governing the concentration behavior. Standard design-effect methodology (Kish, 1965; Bretherton et al., 1999) parameterizes this through an effective sample size $k_{\text{eff}}$ defined such that the variance of the correlated sum equals that of $k_{\text{eff}}$ independent observations:
$\text{Var}(Y)=k_{\text{eff}}\cdot p(1-p).$ (17)
$k_{\mathrm{eff}}\approx\frac{A}{2\pi\xi^{2}},$ (18)
where $\xi$ is the correlation length (Bretherton et al., 1999). Similarly, for temporal correlation with correlation time $\tau$ and unit sampling interval, the Bartlett-Wilks formula (Wilks, 2019) gives:
$k_{\mathrm{eff}}\approx\frac{k}{1+2\sum_{h=1}^{k-1}\rho_{\text{time}}(h)(1-h/k)}.$ (19)
For exponential temporal correlation $\rho_{\text{time}}(h)=e^{-h/\tau}$ with $\tau\gg 1$:
6.3 Impact on Critical Populations
We model the effect of correlation by shrinking the effective number of independent comparisons from $k$ to $k_{\text{eff}}$, holding the marginal match probability $p$ fixed. The true mean number of matches remains $\lambda=kp$, but the concentration (and hence the large-deviation exponent) behaves as if only $k_{\text{eff}}$ coordinates contributed independently.
$q_{\text{corr}}\gtrsim\exp\left(-k_{\text{eff}}p\cdot D(c\|1)\right)=\exp\left(-\lambda\cdot\frac{k_{\text{eff}}}{k}\cdot D(c\|1)\right).$ (21)
(This expression should be interpreted as a heuristic design–effect approximation: positive dependence reduces the large–deviation rate, and replacing $k$ by $k_{\mathrm{eff}}$ captures the dominant variance–inflation effect, but does not constitute a full large–deviation principle under general dependence.)
This heuristic approximation assumes that variance inflation captures the dominant effect of correlation on tail probabilities. Rigorous large-deviation analysis under general dependence is more complex (Dembo and Zeitouni, 2010), but the effective dimension approach provides useful quantitative guidance for practitioners.
This modifies the critical population scale. The exponent in the independent case $\lambda D(c\|1)=kpD(c\|1)$ is reduced by the factor $k_{\text{eff}}/k$, yielding:
$n_{\text{crit}}^{\text{corr}}\lesssim\sqrt{\lambda}\exp\left(\frac{k_{\text{eff}}}{k}\cdot\lambda D(c\|1)\right)=\sqrt{\lambda}\exp\left(k_{\text{eff}}p\cdot D(c\|1)\right).$ (22)
The critical population is thus reduced compared to the independent case, making systems fail at smaller populations when positive correlations are present. The factor $k_{\text{eff}}/k$ represents the information loss due to redundancy in correlated observations.
6.4 Quantitative Examples
Example 6.1 (City-Scale Spatial Surveillance).
$k_{\text{eff}}\approx\frac{100\times 10^{6}\text{ m}^{2}}{2\pi(500)^{2}}\approx 64.$
With $p=0.005$, the expected number of matches is $\lambda=kp=10{,}000\times 0.005=50$. Set the threshold at $m=75$ matches (corresponding to $c=m/\lambda=1.5$, validly in the upper tail regime with $c>1$ as required by Theorem 2.4).
The rate function is:
$D(1.5\|1)=1.5\log(1.5)-1.5+1\approx 0.108.$
$\lambda D(c\|1)=50\times 0.108=5.4,$
giving $n_{\text{crit}}\sim\sqrt{\lambda}\exp(5.4)\approx 7.07\times 221\approx 1{,}560$.
Correlated case: By equation (22), the exponent is reduced by the factor $k_{\text{eff}}/k$:
$\frac{k_{\text{eff}}}{k}\cdot\lambda D(c\|1)=\frac{64}{10{,}000}\times 5.4\approx 0.0345,$
giving $n_{\text{crit}}^{\text{corr}}\sim\sqrt{\lambda}\exp(0.0345)\approx 7.07\times 1.035\approx 7$.
Spatial correlation reduces the critical population from approximately 1,560 to approximately 7, a reduction of over two orders of magnitude. The system can reliably monitor only a handful of individuals once spatial dependencies are properly accounted for, demonstrating how correlation-induced overdispersion dramatically accelerates false-alert saturation.
Example 6.2 (Routine Behavior Monitoring).
For daily surveillance data ($\Delta t=1$ day) over one year ($k=365$ observations) monitoring movement patterns with correlation time $\tau=30$ days, we obtain:
$k_{\text{eff}}\approx\frac{k}{2\tau}=\frac{365}{60}\approx 6.$
With per-observation match probability $p=0.02$, the expected matches are $\lambda=365\times 0.02=7.3$. Set threshold $m=12$ (giving $c=12/7.3\approx 1.64$). The rate function is:
$D(1.64\|1)=1.64\log(1.64)-1.64+1\approx 0.173.$
Independent case: $\lambda D(c\|1)=7.3\times 0.173\approx 1.26$, giving $n_{\text{crit}}\sim\sqrt{\lambda}\,e^{1.26}\approx 2.7\times 3.5\approx 9.5$ (ignoring the subexponential $\sqrt{\lambda}$ factor for rough scaling gives $n_{\text{crit}}\sim e^{1.26}\approx 3.5$).
Correlated case: $(k_{\text{eff}}/k)\cdot\lambda D=(6/365)\times 1.26\approx 0.021$, giving $n_{\text{crit}}^{\text{corr}}\sim\sqrt{\lambda}\,e^{0.021}\approx 2.7\times 1.02\approx 2.8$ (or $e^{0.021}\approx 1.02$ ignoring the prefactor).
Temporal correlation reduces the critical population from approximately 10 individuals to approximately 3, making reliable surveillance of even small groups challenging. Routine behaviors create such massive information redundancy that even year-long observation provides minimal discriminative power.
6.5 Connections to Multiple Testing
This effective dimension reduction connects surveillance analysis to multiple-testing corrections in genomics (Nyholt, 2004) and neuroimaging (Worsley et al., 1996), where similar correlation structures require Bonferroni-type corrections scaled by $k_{\text{eff}}$ rather than the nominal number of tests $k$. The mathematics forces system designers to confront the limited information content of nominally “big” data: more observations do not guarantee more information when those observations are highly correlated.
Remark 6.3 (Overdispersion as System Weakness).
Correlation-induced overdispersion represents a fundamental vulnerability in screening systems. Positive correlation makes extreme events (high match counts) more probable than the independent Poisson model predicts, accelerating the onset of false-alert saturation. System designers cannot escape this by collecting more data; additional correlated observations provide diminishing marginal information while accumulating the same false-positive burden. While our analysis uses effective degrees of freedom as a heuristic to quantify correlation effects, the qualitative conclusion is robust: positive correlation always weakens concentration, making false alerts more likely than independence would suggest.
7 Conclusion

### v2

@@ 4.2 Connections to Algorithmic Fairness
4.3 Limitations and Future Directions
This analysis operates under several simplifying assumptions that define its scope and suggest natural directions for future work.
Independence across individuals. Our core results (Theorems 3.5 and 3.9) assume statistical independence of match counts across individuals (Remark 3.2). Common-mode events (mass gatherings, natural disasters, viral social media content, coordinated activities) introduce positive dependence that would increase false alert rates beyond our bounds. Positive dependence inflates upper-tail probabilities and therefore worsens system reliability relative to the independent case. Extensions via positively associated random variables or Chen–Stein methods [26] could quantify these effects, but our independence-based analysis provides a lower bound on false alert rates.
Binary attributes. We model attributes as binary indicators (match/no-match). Continuous features, count data, or multi-level categorical attributes would require different distributional assumptions and tail bounds. The qualitative insights about combinatorial explosion in high-dimensional spaces should persist, but quantitative thresholds would differ. Extensions to Gaussian or sub-Gaussian attributes would preserve the essential exponential tail behavior underlying our critical-scale results.
Fixed thresholds. Our analysis assumes detection thresholds $m$ are fixed at deployment time. Adaptive systems that adjust thresholds based on observed alert rates or estimated base rates could potentially extend operational lifetimes. However, Theorem 3.9 suggests fundamental limits: under exponential data growth, even optimally adaptive thresholds would need to grow exponentially to maintain reliability, eventually exceeding meaningful detection capabilities.
Constant sensitivity. The Bayesian analysis (Section 4.1) initially treats sensitivity $s=\Pr(\text{flag}\mid\text{target})$ as constant. Remark 4.1 introduces a simple threshold-dependent model, but a complete analysis would require specifying both the target match distribution and the full signal distribution of true targets and performing ROC optimization [23]. This is inherently application-dependent and beyond our current scope.
Heuristic correlation treatment. The effective degrees of freedom approach (Section 3.4; Appendix B) captures variance inflation but does not constitute rigorous large-deviation analysis. Formal treatment would require specifying mixing conditions or dependency graph structures [29, 26]. Our heuristic provides qualitative guidance rather than formal guarantees.
Lack of empirical validation. We have not validated our predictions against operational surveillance data, which is typically proprietary, classified, or subject to confidentiality restrictions. Instead, we use proxy datasets for illustrative rather than validating analyses (Appendix C); these provide qualitative checks but do not constitute validation in the intended deployment environment.
Static population structure. We assume fixed population composition with stable group sizes $n_{g}$ and exposure rates $p_{g}$. Dynamic populations with entry, exit, demographic shifts, and changing surveillance intensity would require stochastic population models. The temporal analysis (Section 3.2) addresses data growth but not population dynamics.
Formal versus heuristic results. Theorems 3.5, 3.9, and 3.13, along with Propositions 3.6 and 3.15, are formal results with complete proofs. The effective dimensionality analysis (Appendix B) and sensitivity-threshold model (Remark 4.1) are heuristic approximations.
Broader applicability. While surveillance systems motivated this analysis, the mathematical framework applies to any domain where threshold rules screen large collections across many low-probability binary indicators. The critical population bounds (Theorem 3.5), temporal saturation dynamics (Theorem 3.9), and group-level disparity amplification (Theorem 3.13) characterize generic properties of high-dimensional threshold detection, independent of the specific application. Natural extensions include network intrusion detection, manufacturing quality control, financial fraud screening, medical diagnostic panels, and environmental monitoring systems. The binary indicator assumption could be relaxed to accommodate hybrid frameworks combining discrete and continuous variables [40], though the essential combinatorial explosion in threshold-based screening would persist. We developed the theory through the surveillance lens because it offered the clearest exposition of the societal stakes, but the probabilistic limits derived here constrain any system that aggregates rare coincidences across high-dimensional attribute spaces.
Despite these limitations, the core mathematical structure (exponential scaling of critical populations, finite system lifetimes under data growth, and structural amplification of exposure disparities) should prove robust across modeling variations. Modeling refinements would shift numerical thresholds but not the qualitative scaling laws, which arise from intrinsic high-dimensional coincidence phenomena.
5 Conclusions